OpenStudy (onepieceftw):

Can someone please help me with this? Find all solutions to the equation. 7 sin^2x - 14 sin x + 2 = -5

OpenStudy (welshfella):

7 sin^2 x - 14 sin x + 7 = 0 this is a quadratic equation which can be solved by factoring

OpenStudy (onepieceftw):

Oh okay, so factored it's:\[(\sin x - 7)(\sin x - 7)\] Now I just need to find the zeros of sinx - 7 = 0